# Scalable approximation of the transformation-free linear simplicial-simplicial regression via constrained iterative reweighted least squares

### v3

@@ Abstract
Simplicial-simplicial regression refers to the regression setting where both the responses and predictor variables lie within the simplex space, i.e. they are compositional. fiksel2022 proposed a transformation-free linear regression model, that minimizes the Kullback-Leibler divergence from the observed to the fitted compositions, where the EM algorithm is used to estimate the regression coefficients. We formulate the model as a constrained logistic regression, in the spirit of tsagris2025, and we estimate the regression coefficients using constrained iteratively reweighted least squares. The simulation studies depict that this algorithm makes the estimation procedure significantly faster, and approximates accurately enough the solution of the EM algorithm.
@@ 1 Introduction
Compositional data [^1] In the field of econometrics they are known as multivariate fractional data (mullahy2015, murteira2016). are non-negative multivariate vectors whose variables (typically called components) convey only relative information. When the vectors are scaled to sum to 1, their sample space is the standard simplex
$\mathbb{S}^{D-1}=\left\{(y_{1},...,y_{D})^{\top}\bigg|y_{i}\geq 0,\sum_{i=1}^{D}y_{i}=1\right\},$ (1)
@@
Examples of such data may be found in many different fields of study and the extensive scientific literature that has been published on the proper analysis of this type of data is indicative of its prevalence in real-life applications [^2] For a substantial number of specific examples of applications involving compositional data see tsagris2020..
The widespread occurrence of this type of data in numerous scientific fields that involve predictors has necessitated the need for valid regression models which in turn has led to several developments in this area, many of which have been proposed recently. Most of these regression models have a restricted attention to the case of a simplicial response (simplicial-real regression setting), or a simplicial predictor (real-simplicial regression setting). The case of simplicial-simplicial regression, where both sides of the equation contain compositional data has not gained too much attention, and this is the main focus of this paper.
@@
Examples with both simplicial responses and predictor variables include wang2013 who modelled the linear relationship between the economic output and inputs in China, and dimarzio2015 who used the composition of the O-horizon layer to predict the composition of the moss layer. chen2017 modelled the relationship between age structure and consumption structure across economic areas. filzmoser2018 who related the percentage of males and females with different education levels across countries. ait2003, alenazi2019 examined how different methods for estimating the composition of white blood cell types are related. chen2021 examined how the the chemical metabolites of Astragali Radix and the plasma metabolites of rats after giving Astragali Radix are related. tsagris2025 studied the relationship between the crop production and the crop cultivated area in Greece and the relationship between voting proportions between successive years in Spain. Finally, rios2025 investigated the association between two high-dimensional multi-omics compositions.
@@
Most published papers regarding simplicial-simplicial regression involve transformations of both simplicial sides. hron2012, wang2013, chen2017 and han2022 used a log-ratio transformation for both the response and predictor variables and performed a multivariate linear regression model. alenazi2019 transformed the simplicial predictors using the $\alpha$-transformation (tsagris2011) followed by principal component analysis and then employed the Kullback-Leibler divergence (KLD) regression (or multinomial logit) model (murteira2016). fiksel2022 who proposed a transformation-free linear regression (TFLR) model whose coefficients lie within the simplex and are estimated via minimization of the Kullback-Leibler divergence (KLD) between the observed and the fitted simplicial responses. The Expectation-Maximisation (EM) algorithm was employed for this purpose.
@@
An issue with the EM algorithm adopted for the estimation of the TFLR’s coefficients is that the EM is in generally computationally slow. To mitigate this computational cost we employ the constrained iteratively reweighted least squares (CIRLS) algorithm, with simplicial constraints imposed on the regression coefficients. Tsagris’s implementation of the EM algorithm (tsagris2025) was shown to be 4 times faster than the implementation of fiksel2022. Our simulation studies demonstrate that CIRLS yields a 2-255 times, depending on the case, speed improvement over the implementation of tsagris2025. The obtained solution is an accurate approximation of the solution produced by the EM algortihm. The speed improvements are particularly important in the following cases: a) having to analyze multiple datasets, b) performing simulation studies, c) working with high dimensional compositions (rios2025), and d) performing permutation or bootstrap based hypothesis tests (fiksel2022, tsagris2025).
The next section describes the TFLR model and the EM algorithm employed to estimate the regression parameters and Section 3 delineates the CIRLS for logistic regression and adopts it to the TFLR model. Section 4 presents the computational complexity of the both algorithms, Section 5 contains simulation studies, followed by the Conclusions section closing the paper.
@@ 2 The TFLR model
fiksel2022 developed the TFLR model, a novel approach for modeling compositional data, addressing methodological limitations inherent in existing methodologies.
The TFLR model is given by:
$\mathbb{E}[\bm{Y}|\bm{X}]=\bm{XB},$ (2)
@@
where $\bm{X}\in\mathcal{S}^{D_{p}-1}$ is a $D_{p}$-dimensional compositional predictor, $\bm{Y}\in\mathcal{S}^{D_{r}-1}$ is a $D_{r}$-dimensional compositional response. The TFLR model relates the $k$-th simplicial response vector to the simplicial predictor via the linear transformation
$E(\bm{Y}_{k}|\bm{X})=\sum_{j=1}^{D_{p}}\bm{X}_{j}B_{jk},$ (3)
where $\bm{B}$ itself belongs to the simplex, $\bm{B}\in\mathbb{R}^{D_{p}\times D_{r}}|B_{jk}\geq 0,\sum_{k=1}^{D_{r}}B_{jk}=1$. Estimation of the elements of $\bm{B}$ takes place by minimizing the KLD of the observed to the fitted compositional vectors
$\min\left\{\sum_{i=1}^{n}\sum_{k=1}^{Dr}y_{ik}\log\left(\frac{y_{ik}}{\sum_{j=1}^{D_{p}}x_{ij}B_{jk}}\right)\right\}=\max\left\{\sum_{i=1}^{n}\sum_{k=1}^{Dr}y_{ik}\log\left(\sum_{j=1}^{D_{p}}x_{ij}B_{jk}\right)\right\}.$ (4)
Unlike conventional approaches that rely on log-ratio transformations (ALR, CLR, ILR), the TFLR model works directly in the simplex space, avoiding transformation-induced interpretability issues. The estimation method, described below, is based on first-moment specifications using estimating equations, avoiding full likelihood specifications and providing robustness to different data-generating mechanisms. The approach naturally accommodates zeros and ones in both predictors and outcomes, addressing a major limitation of log-ratio based methods that require data pre-processing. However, tsagris2025 highlighted that the aforementioned approach allows for zero values in the response variable, but not for the simplicial predictor, in general. Consider for example the case of an observation $\bm{x}_{i}=(x_{i1},x_{i2},0,0)$, where $x_{i1}$ and $x_{i2}$ are different from zero and and the $k$-th column of the estimated matrix $\bm{B}$ is equal to $(0,0,B_{3k},B_{4k})^{\top}$, where both $B_{3k}$ and $B_{4k}$ are not zero. Another possibility is when a column of the $\bm{B}$ matrix is full of zeros. These two cases will produce $x_{ik}B_{ik}=0$ and hence (4) cannot be computed. fiksel2022 circumvents this limitation by adding a small quantity ($\delta=10^{-8}$) during computation of the KLD.
2.1 Expectation-maximisation algorithm
@@
fiksel2022 employed an expectation-maximisation (EM) algorithm to estimate the TFLR model parameters, particularly for the column-stochastic transition matrix $\bm{B}$. They framed the estimation problem as a missing data problem, where each observed compositional vector $\bm{y}_{i}$ is viewed as the result of a two-step generative process. a) For each component $j$ of $\bm{x}_{i}$, a component $k$ of $\bm{y}_{i}$ is selected with probability $b_{jk}$. b) The contributions from all components of $\bm{x}_{i}$ are aggregated according to their proportions $x_{ij}$. The steps of the EM algorithm are described below.
@@
The EM updates naturally satisfy the non-negativity and sum-to-one constraints on each row of $\bm{B}$. Each iteration of the algorithm is guaranteed to increase the likelihood (or pseudo-likelihood) function. The algorithm handles numerical underflow/overflow issues by working with log-transformed probabilities when necessary and accommodates zeros in the data naturally (particularly in the simplicial response), without requiring zero value imputation.
Further, the initial values in the Initialization Step are provided by the simplicially constrained least squares model of tsagris2025. This resulted in an implementation that was shown to be around 4 times faster (tsagris2025) than the implementation in the R package codalm (codalm2021), at the cost of adding extra memory requirements.
@@ 3 TFLR with CIRLS
We first present preliminary concepts regarding the TFLR, the IRLS, the CIRLS and how the latter can be used to estimate the regression coefficients of the TFLR.
@@ 3.2 Logistic regression with identity link and simplicial constraints
Suppose we wish to maximize (5), using the identity link, $\mathbb{E}[y_{i}]=\bm{x}_{i}^{\top}\bm{\beta}$, subject to the following simplex constraints: $0\leq\beta_{j}\leq 1$, for all $j=1,2,\ldots,p$ and $\sum_{j=1}^{p}\beta_{j}=1$. Note, that in this formulation, we have assumed there is no constant. Two notable differences occur, the working response $\bm{z}$ equals the original response vector $\bm{y}$, $z_{i}=y_{i}$ and $w_{i}=\left[(p_{i}(1-p_{i})\right]^{-1}$.
This approach is used to estimate predators diet the quantitative fatty acid signature analysis framework (iverson2004). The standard IRLS algorithm must be modified to incorporate the simplex constraints. At each iteration, after computing the unconstrained update, we need to project the solution back onto the feasible region defined by the simplex. However, this projection approach may not be optimal for likelihood-based inference. Instead, we can formulate the constrained weighted least squares problem at each IRLS iteration as a quadratic programming (QP) problem, yielding the constrained IRLS (CIRLS) formulation (masselot2025).
@@
The QP approach offers several advantages over other methods for handling simplex constraints. The constraints are explicitly included in the optimization problem at each iteration, rather than being enforced after an unconstrained update. QP solvers typically use specialized algorithms that are designed to handle constraints effectively and maintain numerical stability. For convex problems (such as the weighted least squares problem at each IRLS iteration), QP solvers guarantee finding the global optimum subject to the constraints. Unlike approaches that use transformations like the softmax function, QP directly works with the original parameters, avoiding potential numerical issues with transformations. Modern QP solvers are highly optimized and can solve large-scale problems efficiently.
@@
This assigns equal weight to each coefficient. Alternatively, the starting values may come from fitting a constrained OLS with simplicial constraints imposed on the regression coefficients. The convergence criterion we employed was the absolute difference in two successive values of the log-likelihood. In some cases, the optimal solution may lie on the boundary of the simplex, with some coefficients exactly equal to 0 or 1. This can lead to degeneracy in the optimization problem. QP solvers are generally designed to handle such cases, but it may be advantageous to use a solver that employs an interior point method or another approach that handles degeneracy well.
@@ 3.3 CIRLS formulation of the TFLR
To formulate the TFLR as a univariate logistic regression model we follow tsagris2025 and apply the same approach. They constructed the $\bm{D}$ matrix is a $D_{rp}\times D_{rp}$ diagonal matrix, where $D_{rp}=D_{r}\times D_{p}$, and is related to the $\bm{X}^{\top}\bm{X}$ in the following manner
@@
Effectively, we do not vectorise the simplicial responses, or the compositional predictors, and evidently neither the weighting matrix $\bm{W}$. We obtain the $\bm{B}$ matrix from the QP, and subsequently we compute the matrix with the fitted responses, and then the KLD value. We then update the weighting matrix and this process iterates until convergence of the KLD. The initial matrix $\bm{B}$ is obtained from the constrained least squares method of tsagris2025.
The matrix $\bm{D}$ (12) and the vector $\bm{d}$ (13) increase the memory requirements. However, there is no option to reduce the memory requirements and also parallelize the QP step, an important consideration for high-dimensional compositional data.
TFLR solutions may lie on the boundary of the simplex, with some coefficients attaining exactly 0 or $1$. In such cases the active-set approach of (goldfarb1983) is particularly well suited because it handles degeneracy by maintaining a working set of active constraints whose feasibility is enforced at every iteration. This property ensures that the IRLS update always remains in the simplex and that convergence is not impeded by rank deficiencies in the reduced Hessian.
We employ the dual active-set implementation of the R package quadprog (quadprog2019), which is efficient when the number of active constraints at the optimum is small—an empirically common scenario when TFLR solutions lie near the interior of the simplex. Alternative interior-point methods could be used, especially in very high-dimensional regimes where the $O(D_{p}^{3})$ factorisation cost becomes dominant, but we found the active-set method to be numerically stable and highly effective for the dimensionalities considered in this work.

### v4

@@ Abstract
Simplicial–simplicial regression concerns statistical modeling scenarios in which both the predictors and the responses are vectors constrained to lie on the simplex. fiksel2022 introduced a transformation-free linear regression framework for this setting, wherein the regression coefficients are estimated by minimizing the Kullback–Leibler divergence between the observed and fitted compositions, using an expectation–maximization (EM) algorithm for optimization. In this work, we reformulate the problem as a constrained logistic regression model, in line with the methodological perspective of tsagris2025, and we obtain parameter estimates via constrained iteratively reweighted least squares. Simulation results indicate that the proposed procedure substantially improves computational efficiency—yielding speed gains ranging from $6\times--326\times$—while providing estimates that closely approximate those obtained from the EM-based approach.
@@ 1 Introduction
Simplicial, or compositional data, data [^1] In econometrics, these are commonly referred to as multivariate fractional data (mullahy2015, murteira2016). are non-negative multivariate vectors whose components convey only relative information. When the vectors are scaled to sum to 1, their sample space is the standard simplex,
$\mathbb{S}^{D-1}=\left\{(y_{1},\ldots,y_{D})^{\top}\bigg|y_{i}\geq 0,\sum_{i=1}^{D}y_{i}=1\right\},$ (1)
@@
Such data arise in a wide range of scientific disciplines, and the extensive literature on their proper statistical treatment attests to their prevalence in practical applications. [^2] For numerous examples of real-world applications involving simplicial data see tsagris2020.
The frequent appearance of simplicial variables in regression settings has motivated the development of appropriate regression methodologies, leading to several recent methodological advances. Most contributions in this field focus either on the case of a simplicial response with real-valued predictors (simplicial–real regression) or on the converse setting with simplicial predictors and real-valued responses (real–simplicial regression). By contrast, the simplicial–simplicial regression setting, in which both the response and predictor variables are simplicial, has received comparatively limited attention. This constitutes the main focus of the present work.
@@
Applications involving simplicial responses and simplicial predictors include, among others, wang2013, who modeled the relationship between economic outputs and inputs in China, and dimarzio2015, who predicted the composition of the moss layer using the composition of the O-horizon layer. chen2017 investigated the association between age structure and consumption structure across economic regions. filzmoser2018 studied differences in educational composition across countries. ait2003 and alenazi2019 examined relationships between alternative methods of estimating white blood cell type compositions. chen2021 explored associations between chemical metabolites of Astragali Radix and plasma metabolites in rats following administration. tsagris2025 analyzed relationships between crop production and cultivated area in Greece, as well as voting proportion dynamics in Spain. Finally, rios2025 investigated associations between high-dimensional multi-omics simplicial datasets.
@@
Most existing approaches to simplicial–simplicial regression apply transformations to both simplicial variables. For example, hron2012, wang2013, chen2017, and han2022 employed log-ratio transformations for both predictors and responses followed by multivariate linear regression. alenazi2019 transformed simplicial predictors via the $\alpha$-transformation (tsagris2011), applied principal component analysis, and subsequently fitted a Kullback–Leibler divergence (KLD) regression (multinomial logit) model (murteira2016). In contrast, fiksel2022 proposed a transformation-free linear regression (TFLR) framework, in which the regression coefficients lie on the simplex and are estimated by minimizing the KLD between observed and fitted simplicial responses using an Expectation–Maximization (EM) algorithm.
@@
A limitation of the EM-based estimation procedure for the TFLR model is its computational burden, as the EM algorithm is generally slow to converge. To alleviate this issue, we employ a constrained iteratively reweighted least squares (CIRLS) algorithm, incorporating simplex constraints on the regression coefficients. The implementation of the EM algorithm in tsagris2025 was shown to be approximately four times faster than that of fiksel2022. Our simulation studies demonstrate that CIRLS achieves additional substantial computational gains, yielding speed improvements ranging from $2\times$ to $255\times$, depending on the scenario, while producing estimates that closely approximate those obtained via the EM algorithm. These computational benefits are particularly relevant in contexts involving: (a) analysis of multiple datasets, (b) large-scale simulation studies, (c) high-dimensional compositions (rios2025), and (d) permutation- or bootstrap-based inference procedures (fiksel2022, tsagris2025).
The next section presents the TFLR model and the EM algorithm used for parameter estimation. Section 3 describes the CIRLS algorithm for logistic regression and its adaptation to the TFLR framework. Section 4 discusses the computational complexity of both approaches, Section 5 reports simulation results, and the final section provides concluding remarks.
@@ 2 The TFLR model
fiksel2022 developed the TFLR model as a novel framework for modeling simplicial data, addressing methodological limitations inherent in existing approaches.
The TFLR model is defined as
$\mathbb{E}[\bm{Y}|\bm{X}]=\bm{X}\bm{B},$ (2)
@@
where $\bm{X}\in\mathcal{S}^{D_{p}-1}$ denotes a $D_{p}$-dimensional simplicial predictor and $\bm{Y}\in\mathcal{S}^{D_{r}-1}$ denotes a $D_{r}$-dimensional simplicial response. The model relates the $k$-th component of the simplicial response to the predictor through the linear mapping
$E(\bm{Y}_{k}\mid\bm{X})=\sum_{j=1}^{D_{p}}\bm{X}_{j}B_{jk},$ (3)
where the coefficient matrix $\bm{B}$ itself lies in a product of simplices,
$\bm{B}\in\left\{\mathbb{R}^{D_{p}\times D_{r}}\mid B_{jk}\geq 0,\;\sum_{k=1}^{D_{r}}B_{jk}=1\right\}.$
The elements of $\bm{B}$ are estimated by minimizing the Kullback–Leibler divergence between the observed and fitted response compositions:
$\min_{\bm{B}}\left\{\sum_{i=1}^{n}\sum_{k=1}^{D_{r}}y_{ik}\log\left(\frac{y_{ik}}{\sum_{j=1}^{D_{p}}x_{ij}B_{jk}}\right)\right\}\;=\;\max_{\bm{B}}\left\{\sum_{i=1}^{n}\sum_{k=1}^{D_{r}}y_{ik}\log\left(\sum_{j=1}^{D_{p}}x_{ij}B_{jk}\right)\right\}.$ (4)
Unlike conventional approaches relying on log-ratio transformations (ALR, CLR, ILR), the TFLR model operates directly in the simplex, thereby avoiding interpretability challenges and distributional distortions introduced by transformations. The associated estimation procedure is based on first-moment specifications via estimating equations rather than full likelihood specification, providing robustness to a variety of data-generating mechanisms. Furthermore, the method naturally accommodates zeros and ones in both predictors and responses, overcoming a major limitation of log-ratio-based techniques, which typically require preprocessing such as zero-replacement.
However, as noted by tsagris2025, this methodology generally allows for zeros in the response but does not fully accommodate zeros in the simplicial predictors. Consider, for example, an observation $\bm{x}_{i}=(x_{i1},x_{i2},0,0)$, where $x_{i1}$ and $x_{i2}$ are positive, and suppose that the $k$-th column of the estimated matrix $\bm{B}$ is $(0,0,B_{3k},B_{4k})^{\top}$ with $B_{3k}$ and $B_{4k}$ strictly positive. Alternatively, a column of $\bm{B}$ may consist entirely of zeros. In both situations, the product $x_{ik}B_{ik}=0$, rendering (4) undefined. To circumvent this issue, fiksel2022 introduce a small positive constant ($\delta=10^{-8}$) when computing the KLD.
2.1 EM algorithm
@@
fiksel2022 employed the EM algorithm to estimate the parameters of the TFLR model, and in particular the column-stochastic transition matrix $\bm{B}$. The estimation problem is formulated as a missing-data problem, wherein each observed simplicial vector $\bm{y}_{i}$ is interpreted as the outcome of a two-stage generative mechanism. Specifically: (a) for each component $j$ of $\bm{x}_{i}$, a component $k$ of $\bm{y}_{i}$ is selected with probability $b_{jk}$; and (b) the contributions from all components of $\bm{x}_{i}$ are subsequently aggregated according to their proportions $x_{ij}$. The steps of the EM algorithm are outlined below.
@@
The EM updates naturally satisfy the non-negativity and sum-to-one constraints on each row of $\bm{B}$. Moreover, each iteration of the algorithm is guaranteed to increase the (pseudo-)likelihood function. Numerical underflow and overflow issues are mitigated by operating on log-transformed probabilities when necessary, and the algorithm accommodates zeros in the data (particularly in the simplicial response) without requiring zero-value imputation.
In addition, the Initialization Step employs starting values derived from the simplicially constrained least squares estimator of tsagris2025. This enhancement yielded an implementation that was reported to be approximately four times faster than that of the R package codalm (codalm2021), albeit at the cost of increased memory requirements.
@@ 3 TFLR with CIRLS
While the EM algorithm provides a convenient latent-allocation interpretation of the TFLR model, it is not essential for estimating the coefficient matrix $\bm{B}$. The EM procedure maximizes the KLD objective in (4), up to an additive constant, indirectly by introducing latent variables and iteratively improving a lower bound of the KLD criterion.
An alternative perspective arises by noting that the fitted compositions
$\hat{y}_{ik}(\bm{B})=\sum_{j}x_{ij}B_{jk}$
behave as probabilities. Consequently, the optimization problem can be viewed as a constrained regression problem defined on the simplex. This viewpoint motivates the use of iteratively reweighted least squares (IRLS), which replaces the EM lower bound with a quadratic surrogate obtained via a second-order Taylor expansion of the objective function.
Thus, both EM and IRLS yield weighted updates for $\bm{B}$, although IRLS generates Newton-type additive steps that typically require substantially fewer iterations. Each IRLS update reduces to solving a strictly convex quadratic program subject to simplicial constraints, leading to a computationally efficient approximation to the EM solution while preserving the structure of the parameter space.
In what follows, we present preliminary concepts regarding the TFLR model, the IRLS procedure, and its constrained extension (CIRLS), and demonstrate how the latter can be applied to estimate the regression coefficients of the TFLR framework.
@@ 3.2 Logistic regression with identity link and simplicial constraints
Suppose we wish to maximize Eq. (5), using the identity link, $\mathbb{E}[y_{i}]=\bm{x}_{i}^{\top}\bm{\beta}$, subject to the following simplex constraints: $0\leq\beta_{j}\leq 1$, for all $j=1,2,\ldots,p$ and $\sum_{j=1}^{p}\beta_{j}=1$. Note, that in this formulation, we have assumed there is no constant. Two notable differences occur, the working response $\bm{z}$ equals the original response vector $\bm{y}$, $z_{i}=y_{i}$ and $w_{i}=\left[(p_{i}(1-p_{i})\right]^{-1}$.
This approach has been used, for example, to estimate predator diet within the quantitative fatty acid signature analysis framework (iverson2004). However, the standard IRLS algorithm must be modified to accommodate simplex constraints. In particular, at each iteration, after computing the unconstrained update, one would need to project the estimate back onto the feasible region defined by the simplex. Such projection-based corrections, however, may not be optimal from a likelihood-based perspective. A more principled alternative is to formulate the constrained weighted least squares problem at each IRLS iteration as a quadratic programming (QP) problem, leading to the constrained IRLS (CIRLS) formulation (masselot2025).
@@
The QP approach offers several advantages over alternative methods for handling simplex constraints. First, the constraints are explicitly incorporated into the optimization problem at each iteration, rather than being imposed post hoc following an unconstrained update. Moreover, QP solvers employ specialized algorithms designed to efficiently manage constraints while maintaining numerical stability. For convex optimization problems—such as the weighted least squares problem arising at each IRLS iteration—QP solvers guarantee convergence to the global optimum under the imposed constraints. In contrast to transformation-based approaches (e.g., softmax reparametrization), QP directly operates on the original parameters, thereby avoiding potential numerical difficulties associated with reparametrization. Finally, modern QP solvers are highly optimized and capable of handling large-scale problems with considerable computational efficiency.
@@
This approach assigns uniform weights to each regression coefficient. As an alternative initialization strategy, starting values can be obtained by fitting ordinary least squares under simplicial constraints on the coefficient vector. We employ an absolute difference convergence criterion based on successive log-likelihood values. Under certain parametric configurations, the optimal solution may reside on the boundary of the simplex, resulting in one or more coefficients attaining their boundary values of 0 or 1. Such boundary solutions can induce degeneracy in the optimization problem. While quadratic programming solvers are typically equipped to handle degenerate cases, implementation with solvers employing interior point methods or alternative algorithms designed for robust handling of degeneracy may prove more reliable and computationally stable.
@@ 3.3 CIRLS formulation of the TFLR
To formulate the TFLR as a univariate logistic regression model we follow tsagris2025. They constructed the $\bm{D}$ matrix is a $D_{rp}\times D_{rp}$ diagonal matrix, where $D_{rp}=D_{r}\times D_{p}$, and is related to the $\bm{X}^{\top}\bm{X}$ in the following manner
@@
Notably, we do not vectorize the simplicial responses or the simplicial predictors, nor the weighting matrix $\bm{W}$. The matrix $\bm{B}$ is obtained via quadratic programming, from which we subsequently compute the fitted response matrix and the corresponding KLD value. The weighting matrix is then updated, and this iterative process continues until the KLD converges. Initial values for $\bm{B}$ are obtained using the constrained least squares method of tsagris2025.
The construction of matrix $\bm{D}$ (12) and vector $\bm{d}$ (13) imposes substantial memory requirements. Unfortunately, no straightforward approach exists to simultaneously reduce memory demands and parallelize the quadratic programming step—a critical consideration when analyzing high-dimensional simplicial data.
Solutions obtained via TFLR may reside on the boundary of the simplex, with certain coefficients attaining their limiting values of exactly 0 or $1$. Under such circumstances, the active-set approach of goldfarb1983 proves particularly well suited, as it handles degeneracy by maintaining a working set of active constraints whose feasibility is rigorously enforced at each iteration. This mechanism ensures that IRLS updates remain within the simplex and that convergence is not compromised by rank deficiencies in the reduced Hessian.
We employ the dual active-set implementation provided by the R package quadprog (quadprog2019), which exhibits high computational efficiency when the number of active constraints at the optimum is small—a scenario frequently observed empirically when TFLR solutions lie near the interior of the simplex. While alternative interior-point methods could be deployed, particularly in high-dimensional regimes where the $O(D_{p}^{3})$ factorization cost becomes dominant, we find the active-set method to be both numerically stable and highly effective across the dimensionalities considered in this work.